HUBBLE SPACE TELESCOPE DAILY REPORT #2144 PERIOD COVERED: 0000Z (UTC) 06/12/98 - 0000Z (UTC) 06/15/98 Daily Status Report as of 166/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed Three Sets of STIS/CCD 7926 (CCD Dark and Bias Monitor -- Continued) The Space Telescope Imaging Spectrograph (CCD) was used to make several dark and bias calibration observations. This is to measure CCD dark current and bias rates for the two supported amplifier settings {gain = 1 and gain = 4} in order to gather data for making superdarks and superbiases. These observations will also track the growth of hot pixels on the CCD in general and within the various target acquisition apertures specifically. The observations were executed as scheduled, and no anomalies were noted. 1.2 Completed Twenty Sets of STIS/CCD 7908 (POMS Test Proposal: STIS Non-Scripted Parallel Proposal) The Space Telescope Imaging Spectrograph (CCD) was used to make calibration observations as part of the POMS Test Proposal. This was a STIS non-scripted parallel proposal. During the first iteration of this proposal, the late acquisition and four losses of lock described in HSTAR 6528 and in the previous report occurred, affecting one observation. During the third iteration of this proposal, there was a loss of lock as described in 2.1 (and HSTAR 6532) that affected one observation. During the ninth iteration of this proposal, the acquisition failures mentioned in 2.1 and HSTAR 6536 occurred. During the tenth iteration, the ARU/PRT was uplinked, and from the re-acquisition all observations were successful. During the fifteenth iteration, the fine lock backup mentioned in 2.1 and HSTAR 6539 occurred, possibly affecting the observations. Otherwise, the observations were completed as planned, and no further problems were reported. 1.3 Completed STIS/MA2 7604 (Cycle 7 MAMA Dark Measurements) The Space Telescope Imaging Spectrograph (MA2) was used to make MAMA dark measurements. These darks will be used in later image reductions and also trended over the instrument's life time as an indicator of the detector's degradation or damage. The observations were completed as scheduled, and no problems were reported. 1.4 Completed Four Sets of NIC/3 7872 (A Morphological Census of z > 1 Cluster Galaxies in the Optical Rest-Frame) The Near Infrared Camera (Camera Three) was used to make observations of the galactic cluster containing 3C267 P1, P2 and P3. Our existing WF/PC-2 imaging of this cluster has let us study the rest-frame ultraviolet light from these galaxies. The proposed NICMOS imaging will allow us to morphologically classify the cluster galaxies at rest-frame optical wavelengths and to place them into the context of the Hubble Sequence. During the first iteration of this proposal, the two losses of lock detailed in HSTAR 6530 and in the previous report occurred. During the second iteration of this proposal, there was a loss of lock as described in 2.1 (and HSTAR 6532) that affected one observation. During the third iteration, the losses of lock mentioned in HSTAR 6534 and 2.1 occurred, affecting three observations. Otherwise, the observations were executed as scheduled, and no further anomalies were noted. 1.5 Completed WF/PC-2 8049 (Long WF/PC-2 Decontaminations during 2nd NIC/3 Campaign) The WF/PC-2 was used to make long decontamination procedures as part of the second NIC/3 campaign. Each decon is followed by five 30-minute darks. During this time, no external science WF/PC-2 observations were executed. During this proposal, the losses of lock mentioned in HSTAR 6534 and 2.1 occurred, affecting three observations. Otherwise, the decontaminations were executed as scheduled, and no further anomalies were noted. 1.6 Completed NIC/3 7816 (NIC/3 Photometric Zero Point) The Near Infrared Camera (Camera Three) was used to make observations of standard star BRI0021-CA in order to study the zero point flux for a set of filters. The NIC/3 intra-pixel sensitivity will also be measured. The observations were executed as scheduled, and no anomalies were noted. 1.7 Completed NIC/3 7830 (IR Spectra for Known-Mass M Dwarfs) The Near Infrared Camera (Camera Three) was used to make infrared spectrographic observations of M-Class dwarf star GJ623AB. Over the last decade there has been considerable debate on two major issues in the low-mass star literature: {1} the effective temperature scale for M dwarfs and its relation to mass and luminosity and {2} the conversion from an absolute magnitude in some passband to a bolometric magnitude. The observations were executed as scheduled, and no anomalies were noted. 1.8 Completed Seventeen Sets of NIC/3 7919 (NICMOS Camera 3 Campaign Snapshots) The Near Infrared Camera (Camera Three) was used to make images of galaxies NGC-224, NGC-7309, NGC-5605, NGC-3571, NGC-128, NGC-6754, NGC-672, NGC-4085, NGC-4144, NGC-3379, NGC-4293, NGC-4449, NGC-598, NGC-4389, NGC-2787 and NGC-4183 as part of the NICMOS Camera 3 campaign snapshots. The proposal is to obtain wide field snapshot images of nearby galaxies in Paschen Alpha which is inaccessible from the ground. During the first iteration of this proposal, the acquisition failures mentioned in 2.1 and HSTAR 6536 occurred. During the seventh iteration, the fine lock backup mentioned in 2.1 and HSTAR 6539 occurred, possibly affecting the observations. Otherwise, the observations were executed as planned, and no further anomalies were reported. 1.9 Completed Two Sets of STIS/CCD 7912 (STIS Parallel Archive Proposal - Nearby Galaxies - Imaging and Spectroscopy) The Space Telescope Imaging Spectrograph (CCD) was used to make parallel observations of nearby galaxies. This survey will be useful to study the star formation histories, chemical evolution, and distances to these galaxies. These data will be placed immediately into the Hubble Data Archive. The observations were completed as scheduled, and no problems were reported. 1.10 Completed Four Sets of NIC/1/2/3 7817 (A Complete NICMOS Map of the Hubble Deep Field) The Near Infrared Camera (Cameras One, Two and Three) was used to make observations to map the complete WF/PC-2 region of the central Hubble Deep Field. These data will provide unique information about the faintest objects, but will include only a small number of objects with spectroscopic redshifts or z ~ 3 Lyman break galaxies. During the first and second iterations of this proposal, the re-acquisitions failed that are mentioned in HSTAR 6536 and in 2.1 occurred. During the third iteration, the losses of lock mentioned in 2.1 and HSTAR 6538 occurred. Otherwise, the observations were executed as scheduled, and no further anomalies were noted. 1.11 Completed Four Sets of WF/PC-2 8013 (WF/PC-2 Parallel Images to the HDF Flanking Fields {Companion Proposal to 7817}) The WF/PC-2 was used to make observations parallel to the Hubble Deep Field observations (i.e., flanking fields). The complete WF/PC-2 region of the central Hubble Deep Field {HDF} was being mapped with the NICMOS Camera 3 at F110W and F160W {proposal 7817}. While these observations were executing, the WF/PC-2 was working in parallel mode to image areas near this field. During the first iteration of this proposal, the re-acquisitions failed that are mentioned in HSTAR 6536 and in 2.1 occurred. During the second iteration, the acquisitions failure mentioned in the same HSTAR occurred. The recovery also occurred during this iteration. During the third iteration, the losses of lock mentioned in 2.1 and HSTAR 6538 occurred. Otherwise, the observations were completed as scheduled, and no further anomalies were reported. 1.12 Completed NIC/1/2/3 7239 (Surface Compositions of Kuiper Belt Objects) The Near Infrared Camera (Cameras One, Two, and Three) was used to make observations of the Kuiper Belt object Pholus in order to study its surface composition. The surfaces of Kuiper Belt Objects contain a record of formative and dynamical processes in the outer solar system. The observations were executed as scheduled, and no anomalies were noted. 1.13 Completed NIC/1/2/3 7822 (Grism Spectra of Centaurs and TNOs from 1-2 Microns) The Near Infrared Camera (Cameras One, Two and Three) was used to make grism spectra of object 1997CU26 from 1-2 microns. The surfaces of Centaurs and Trans-Neptunian Objects (TNOs) contain potentially unique records of early evolutionary processing {accretional and dynamical} in the outer solar system. The observations were executed as scheduled, and no anomalies were noted. 1.14 Completed STIS/CCD 8051 (STIS Out-of-Focus PSF) The Space Telescope Imaging Spectrograph (CCD) was used to make observations of standard source GRW+70D582 in order to study the out-of-focus point spread function. The observations were completed as scheduled, and no anomalies were reported. 1.15 Completed NIC/1/2/3 7223 (Spectrophotometry and Imaging of Pluto and Charon) The Near Infrared Camera (Cameras One, Two, and Three) was used to make observations of Charon. Detailed spectra of the individual members of the Pluto-Charon system are not known. The observations were completed as scheduled, and no problems were reported. 1.16 Completed NIC/3 7892 (Narrow-band Imaging of BAL QSOs) The Near Infrared Camera (Camera Three) was used to make narrow band observations of the BL-Lac (BAL) QSO Q1225+1512. The detection of extended emission-line gas structures in BAL QSOs would offer another means of investigating the BAL QSO phenomenon as well as establishing a new probe of material in galaxies at high redshift. The observations were completed as planned, and no problems were reported. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 34 Successful : 32 Scheduled Reacquisitions : 15 Successful : 13 HSTAR 6532 documents a loss of lock at 163/0928Z after a T3 slew. The proposals described in 1.2 1nd 1.4 were affected. HSTAR 6534 documents multiple losses of lock around 163/125745Z, following a profiled (T3) slew. The proposals described in 1.4 and 1.5 were affected. The re-acquisition at 164/0506Z failed while the vehicle was in fine lock. A map at 164/0513Z showed a 127-second arc-second roll error, and an ARU/PRT slew was uplinked at 164/0618Z. This slew was rejected since an FGS clearing slew (not shown on the MEGG timeline) was in progress. Another re-acquistion at 164/063Z also failed. Then at 164/0901Z and at 164/0935Z two primary acquisitions failed. The SPC error flag was reset at 164/0819 per ROP DF-16. Another ARU/PRT was successfully uplinked at 164/0901Z. The following re-acquisition at 164/0935Z was successful. These events are described in HSTAR 6536. The proposals described in 1.2, 1.8, 1.10, and 1.11 were affected. HSTAR 6538 documents multiple losses of lock near 165/0528Z and 165/0705Z following T3 slews. The proposals described in 1.10 and 1.11 were affected. The acquisition at 165/165033Z failed to fine lock backup on FGS-3 only when the FGS-1 scan step limit was exceeded. The proposals described in 1.2 and 1.8 may have been affected. 2.2 FHST Updates: Scheduled : 53 Successful : 52 As described in HSTAR 6533 the roll delay update scheduled at 163/134320Z failed. The subsequent guide star acquisition AT 163/145529Z succeeded. 2.3 Operations Notes: The SSR EDAC error counter was reset four times using ROP SR-1. DF-224 SPC dumps were collected at 164/1544Z (on B String) and at 165/1800Z (on A String) for CCS testing purposes. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science observations and calibrations. /DMH